OpenStudy (anonymous):

Write an algebraic equation for the following. A number raised to the fifth power is four less than ten times the number. A. 5x=10x-4 B. x^5=4-10x C. 5x=4-10x D. x^5=10x-4

OpenStudy (anonymous):

well x^5 is a number raised to the 5th power so set that = to the other part since it says is refers to = sign... then 4 less is -4 + 10 time the number which is x so the equation should be similar to x^5 = -4 + 10x.

OpenStudy (anonymous):

thank you:) could you help me with this last one I can not get it at all. The quotient of a number and four is twelve less than the number. What is the number? @fiberdust

OpenStudy (anonymous):

the quotient i know is a division term so i think it means 1 divided by (x+4) = for is then less than 12 would be - 12 + x so equation woulbe be 1/(x+4) = x - 12 but i might be wrong on the 1 / (x+4).

OpenStudy (anonymous):

looks like it's supposed to be x / 4 = x - 12

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

your welcome i'm glad i could help hope i didn't confuse u any